If the price of profiteroles decreases from $1.10 to $0.90, and the quantity demanded of profiteroles does not change, then the price elasticity of demand, obtained using the midpoint method, is:




1
Expert's answer
2021-10-20T15:55:10-0400

% change in price"=\frac{1.10-0.90}{\frac{1.10+0.90}{2}}\times100=20"%

% change in quantity = 0%.

Price Elasticity of Demand "=\frac{0}{20}=0"


The value of elasticity indicates that demand for the product is perfectly inelastic. This is because a change in price has no impact on quantity demanded.